# Transparent Graphene-Superconductor Interfaces: Quantum Hall and Zero Field Regimes

### v1

@@ IV.3 Spectrum of the CAEM
We now turn to discussing in more detail the energy spectrum of the graphene-superconductor system. Fig. 8 shows an example for each stitching when the superconductor is well-matched and the interface is transparent [89]. In each plot, the BdG spectrum for two values of the superconducting gap are superposed, $\Delta/E_{L}\!=\!0.1$ (green) and $0.5$ (blue). The smaller value is more realistic, but results for the larger value are instructive because the bulk LLL QH states are within the superconducting gap, $\Delta>\mu_{\textrm{gr}}=E_{L}/4$, making the figures clearer. With the gauge that we use, the QH states with guiding center coordinate at the interface, $\overline{y}\!=\!0$, correspond to $k=K\!=\!2\pi/3a$ and $k=K^{\prime}\!=\!4\pi/3a$.
For the larger superconducting gap (blue), the energies of the bulk LLL states are clear (horizontal blue lines at ±$\pm\mu_{\textrm{gr}}$); they mark the energy of the electron and hole Dirac points. The range of the superconductor’s continuum states—the thicket of states at energies $>\Delta$—differs for the three cases due to the size of the Fermi surface. It is largest for dense stitching [panel (a)] in which the superconductor’s band structure is folded back into the graphene Brillouin zone because the period of the square lattice is $a/2$. For the smaller superconducting gap (green), the QH states hybridize with the superconducting continuum and largely disappear as a distinct signature in the spectrum.
Two CAEM are present at the chemical potential near $k=K$ or $K^{\prime}$, while the two terminated-lattice edge states on the other edge of the nanoribbon are basically unchanged [compare to Fig. 6(c)]. The velocity of the CAEM are substantially reduced compared to those at a terminated-lattice edge, especially for the smaller value of $\Delta$ (see insets). This velocity reduction results from the slow decay of the wavefunction in the superconductor which implies that the particles spend a substantial amount of time there. The velocity reduction is consistent with the QH rule of thumb [[][p.311.]IhnBook] that a softer barrier leads to a smaller velocity.
@@
where $\delta k$ is the deviation from the corresponding valley center. Combining this with the particle-hole symmetry relation for a spin-independent BdG equation, $E_{K}(\delta k)=-E_{K^{\prime}}(-\delta k)$, one finds that the dispersion is antisymmetric within each valley,
@@
Hand in hand with these changes in the CAEM dispersion, the presence of a barrier changes the electron-hole hybridization of the state at the chemical potential. Fig. 10 shows the hybridization as a function of the barrier strength in the continuum model, for the case of perfect hybridization when $Z=0$ ($\widetilde{v}^{2}\!=\!1$). The presence of a barrier at the interface prevents contact with the superconductor and thus drives the $K$-valley mode toward being fully electron-like [panel (a)]. The effect of the barrier is more pronounced for lower filling where the particles have less kinetic energy. On the other hand, panel (b) shows that a potential well at the interface, $Z<0$ with $|Z|\sim 1$, can enhance the hole content.
@@
IV.4 Valley Degeneracy, Interface Transparency, and Andreev Reflection
Valley degeneracy is not accidental but rather is a universal feature of well-hybridized states [77]. Moreover, valley degeneracy is independent of the graphene chemical potential (for $\mu_{\textrm{gr}}$ not too close to the Dirac point): the interface mode “self-aligns” so that $E_{K}(\delta k=0)=0$ throughout the LLL. We find that this holds for all of the well-matched, highly transparent tight-binding and continuum cases that we have studied.
@@ IV.4 Valley Degeneracy, Interface Transparency, and Andreev Reflection
In the LLL of the terminated lattice spectrum, Fig. 6, such a state exists only for the fine-tuned case $\mu_{\textrm{gr}}=E_{L}$ [84, 78]. In contrast, for a transparent interface, $t_{NS}\sim t\gg E_{L}\gg\Delta$, an approximately Gaussian wavefunction centered at $\delta\overline{y}_{0}\!=\!0$ can be constructed, by mixing in small amplitudes of higher Landau level edge states. Alternatively, using the bulk QH complete basis, a small modification of the LLL wavefunction at $\delta\overline{y}\!=\!\delta k\!=\!0$ shifts its energy from the Dirac point to the chemical potential. Thus, we see that for a transparent interface a “self-aligned” $E\!=\!0$ interface mode, i.e. $\delta\overline{y}_{0}\!=\!\delta k_{0}\!=\!0$ with no tuning, is natural.
This self-aligned interface mode is evident in the tight-binding numerics, Figs. 5, 8, and 9, as well as the continuum results, Fig. 1. Once $E(\delta k\!=\!0)=0$ is established, full valley degeneracy follows in two steps. First, the spectrum within a given valley should be antisymmetric in $\delta k$ because the role of electrons and holes is exactly conjugate for a transparent interface. Second, the BdG particle-hole symmetry then implies valley degeneracy.
@@ V.0.1 Tight-binding model
The spectra here are qualitatively similar to each other but distinctly different from the optimal case (Fig. 8) in that valley degeneracy is absent. Valley degeneracy implies antisymmetry, Eqs. (24)-(25), and the black lines in the insets [which are $-E_{K}(-\delta k)$] clearly do not overlay the blue and green. Quantitatively, the deviation of the $E\!=\!0$ state from $K$, $\delta k_{0}$, is given in Table 1 for both cases: though much larger than for the well-matched cases, these values are still smaller than for the QH terminated-lattice edge state  [^5] In the continuum model, such a deviation of the propagating mode at the chemical potential, $\delta k_{0}$ is obtained for $\tilde{v}^{2}=0.38$ and $0.41$, respectively.. Because $\delta k_{0}$ is of order $1/l_{B}$, the guiding center coordinates of the electron and hole components differ by of order $l_{B}$, and thus their spatial profiles differ substantially.
@@ V.0.2 Continuum model
Finally, we present in Fig. 15 results when both a detuned $\mu_{S}$ and a small interface barrier, $Z=0.5$, are present. There continues to be only a modest dependence on the Fermi surface of the superconductor through $\widetilde{v}^{2}$, but there is somewhat more dependence on the graphene chemical potential $\mu_{\textrm{gr}}$. Note, however, that the dependence is smooth: there continues to be no indication that fine-tuning is needed to obtain strong Andreev reflection.
@@ VI Spin: Zeeman-Split Modes
Figure 17: Wavefunctions, $|\psi(y)|^{2}$, for the two Zeeman-split modes in the $K$ valley at the chemical potential ($E=0$): (a) spin up and (b) spin down solutions. The sparse-stitching interface is well matched and so maximally transparent. The degree of $e$-$h$ hybridization is given in Table 2. The magenta line is the zero-Zeeman case for comparison. (Correspond to spectrum in Fig. 16(a).)
@@
Figure 18: For an imperfect interface, wavefunctions for the two $E\!=\!0$ modes in the $K$ valley as they are split by the Zeeman effect: (a) spin up and (b) spin down solutions. Here, $\mu_{S}/t=4.2$. The magenta line is the zero-Zeeman case for comparison (not plotted in superconductor for clarity). (Correspond to spectrum in Fig. 16(b).)
@@ VII Conclusions
We have presented results for an infinite graphene/superconductor interface using continuum and tight-binding models in the standard mean-field approach to spatially inhomogeneous superconductivity (BdG equation). Our main assumptions are: (i) disorder is absent and (ii) the fields $\Delta$, $B$, and $U$ are taken to be step functions at the interface. The continuum results are particularly helpful in exploring a wide range of parameters.
The primary conclusion is that strong Andreev reflection, and hence strong electron-hole hybridization, is much easier to achieve with graphene than with a typical semiconductor 2D electron gas. The reason is that the velocity at the Dirac points in graphene is large—of order the Fermi velocity of metals—and independent of the graphene chemical potential or density. Thus, current continuity across the interface can be satisfied without a great deal of electron reflection.
In order to actually realize strong Andreev reflection, the interface should be fairly transparent. In particular, there should be no obvious barrier, and the superconductor should be well-matched, having Fermi level states with wavevector larger than $K$ and velocity of order $v$ (corresponding to the wave vector of the Dirac point and the Fermi velocity of graphene). This does not appear to be a stringent restriction, since the Fermi surface of metals used for thin-film superconductors often have several sheets, allowing for several possible matches.
Previous work has emphasized that fine-tuning is necessary in order to achieve electron-hole hybridization at a clean QH/superconductor interface [45, 57, 58]; for instance, the chemical potential in the QH material should be tuned on an energy scale less than $E_{L}$. We emphasize that this is not the case for graphene. Once one obtains a fairly well-matched, transparent interface, which may involve “coarse-tuning” the choice of superconductor on an energy scale of the bandwidth ($\gg\!E_{L}$), the $e$-$h$ hybridization in a graphene CAEM is ideal and independent of filling. Indeed, the dispersion of the CAEM self-aligns to be antisymmetric within each valley, independent of filling. Thus, the spectrum is valley degenerate, and the guiding center coordinate of the state at the chemical potential is at the graphene/superconductor interface.
@@
In the spatial profile of the wavefunction, the weight of the electron component is primarily on the sublattice corresponding to the bulk LLL state (sublattice-valley locking), and is surprisingly close to a truncated bulk wavefunction. This demonstrates that the terminated-lattice QH edge states are a poor starting point for understanding the CAEM of a transparent interface, an unsurprising result in hindsight in view of the very different boundary conditions.
In future work, the role of disorder, both in the superconductor and along the interface, should be further developed in the context of a transparent graphene/superconductor interface (for initial work see Ref. [53] and for the importance of disorder in the tunneling regime see Ref. [57]). We suspect that disorder will decrease the dependence on $\mu_{S}$ in our models and so increase the window of good electron-hole hybridization. A second clear direction is to include the penetration of the magnetic field into the superconductor in the form of vortices. It is likely that tunneling of electrons into states in the vortices is an important source of loss in the experiments [29, 30], and theoretical work assessing the importance of this effect has recently appeared [55, 60, 63]. Third, the role of spin-orbit coupling should be investigated. While such coupling is weak in graphene, in several of the superconductors typically used it is strong. Finally, perhaps the most straight forward extension of our results is to the properties of nanostructures, which involves scattering between these CAEM at junctions, corners, and quantum point contacts.
@@ Acknowledgements.
We thank Anthony David, Gleb Finkelstein, William Klein, and Lingfei Zhao for helpful discussions. The work in the USA was supported in part by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Materials Sciences and Engineering Division, under Award No. DE-SC0005237. A.B. acknowledges support by the 2019 Department of Defense NDSEG fellowship. G.Z. acknowledges support from the National Natural Science Foundation of China (Grant No. 12374158) and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302400).
@@
Appendix A Specifics of the Models: Tight-Binding and Continuum
@@
$H_{T,\text{dense}}=$ $H_{T}^{(A)}+H_{T}^{(B)}$

### v2

@@ IV.3 Spectrum of the CAEM
We now turn to discussing in more detail the energy spectrum of the graphene-superconductor system. Fig. 8 shows an example for each stitching when the superconductor is well-matched and the interface is transparent [91]. In each plot, the BdG spectrum for two values of the superconducting gap are superposed, $\Delta/E_{L}\!=\!0.1$ (green) and $0.5$ (blue). The smaller value is more realistic, but results for the larger value are instructive because the bulk LLL QH states are within the superconducting gap, $\Delta>\mu_{\textrm{gr}}=E_{L}/4$, making the figures clearer.
For the larger superconducting gap (blue), the bulk LLL states at the energies of the electron and hole Dirac points are clear (horizontal blue lines at ±$\pm\mu_{\textrm{gr}}$). The range of the superconductor’s continuum states—the thicket of states at $E>\Delta$—differs for the three cases due to the size of the Fermi surface. It is largest for dense stitching [panel (a)] in which the superconductor’s band structure is folded back into the graphene Brillouin zone. For the smaller superconducting gap (green), the QH states hybridize with the superconducting continuum and largely disappear as a distinct signature in the spectrum.
Two CAEM are present at the chemical potential near $k\!=\!K$ and $K^{\prime}$, while the two terminated-lattice states on the other edge of the nanoribbon are unchanged [compare to Fig. 6(c)]. The velocity of the CAEM are substantially reduced compared to a terminated-lattice edge, especially for the smaller $\Delta$ (see insets). This velocity reduction results from the slow decay of the wavefunction in the superconductor which implies that the particles spend a substantial amount of time there. The velocity reduction is consistent with the QH rule of thumb [[][p.311.]IhnBook] that a softer barrier leads to a smaller velocity.
@@
where $\delta k$ is the deviation from the corresponding valley center. Combining this with BdG particle-hole symmetry $E_{K}(\delta k)=-E_{K^{\prime}}(-\delta k)$, one finds that the dispersion is antisymmetric within each valley,
@@
Hand in hand with these changes in the CAEM dispersion, the presence of a barrier changes the electron-hole hybridization of the state at the chemical potential. Fig. 10 shows the hybridization as a function of the barrier strength in the continuum model ($Z\!=\!V_{0}/\hbar v$), for the case of perfect hybridization ($\widetilde{v}^{2}\!=\!1$) when $Z\!=\!0$. The presence of a barrier at the interface prevents contact with the superconductor and thus drives the $K$-valley mode toward being fully electron-like [panel (a)]. The effect of the barrier is more pronounced for lower filling where the particles have less kinetic energy. On the other hand, panel (b) shows that a potential well at the interface, $Z<0$ with $|Z|\sim 1$, can enhance the hole content.
@@
IV.4 Valley Degeneracy, Interface Transparency,
and Andreev Reflection
Valley degeneracy is not accidental but rather is a universal feature of well-hybridized states [46]. Moreover, valley degeneracy is independent of the graphene chemical potential (for $\mu_{\textrm{gr}}$ not too close to the Dirac point): the interface mode “self-aligns” so that $E_{K}(\delta k\!=\!0)\!=\!0$ throughout the LLL [shown explicitly for the continuum model in Fig. 1(b)]. We find that this holds for all of the well-matched, highly transparent tight-binding and continuum cases that we have studied.
@@ IV.4 Valley Degeneracy, Interface Transparency,
In the LLL of the terminated lattice spectrum, Fig. 6, no such a state exists. In contrast, for a transparent interface, $t_{NS}\sim t\gg E_{L}\gg\Delta$, an approximately Gaussian wavefunction centered at $\delta\overline{y}_{0}\!=\!0$ can be constructed, by mixing in small amplitudes of higher Landau level edge states. Alternatively, using the bulk QH complete basis, a small modification of the bulk LLL wavefunction at $\delta\overline{y}\!=\!\delta k\!=\!0$ triggered by $t_{NS}$ shifts its energy from the Dirac point to the chemical potential. In this way, a large interface matrix element $t_{NS}$ can produce both an electron and a hole at the chemical potential centered on the interface, which are then coupled by the pairing matrix element $\Delta$. Thus, we see that for a transparent interface a “self-aligned” $e$-$h$ interface mode, i.e. $\delta\overline{y}_{0}\!=\!\delta k_{0}\!=\!0$ with no tuning, is natural.
This self-aligned interface mode is evident in the tight-binding numerics, Figs. 5, 8, and 9 and Table 1, as well as the continuum results, Fig. 1. Once $E(\delta k\!=\!0)=0$ is established, full valley degeneracy follows in two steps. First, the spectrum within a given valley should be antisymmetric in $\delta k$ because the role of electrons and holes is exactly conjugate for a transparent interface. Second, the BdG particle-hole symmetry then implies valley degeneracy.
@@ V.0.1 Tight-binding model
The spectra here are qualitatively similar to each other but distinctly different from the optimal case (Fig. 8) in that valley degeneracy is absent. Valley degeneracy implies antisymmetry, Eqs. (24)-(25), and the black lines in the insets [which are $-E_{K}(-\delta k)$] clearly do not overlay the blue and green. Quantitatively, the deviation $\delta k_{0}$ of the $E\!=\!0$ state from $K$ is given in Table 1 for both cases: though much larger than for the well-matched cases, these values are still smaller than for the QH terminated-lattice edge state  [^5] In the continuum model, such a deviation of the propagating mode at the chemical potential, $\delta k_{0}$ is obtained for $\tilde{v}^{2}=0.38$ and $0.41$, respectively.. Because $\delta k_{0}$ is of order $1/l_{B}$, the guiding center coordinates of the electron and hole components differ by of order $l_{B}$, and thus their spatial profiles differ substantially.
@@ V.0.2 Continuum model
Finally, we present in Fig. 15 results when both a detuned $\mu_{S}$ and a small interface barrier, $Z=0.5$, are present. There continues to be only a modest dependence on the Fermi energy of the superconductor through $\widetilde{v}^{2}$, but there is somewhat more dependence on the graphene chemical potential $\mu_{\textrm{gr}}$. Note, however, that the dependence is smooth: there continues to be no indication that fine-tuning is needed to obtain strong Andreev reflection.
@@ VI Spin: Zeeman-Split Modes
Figure 17: Wavefunctions, $|\psi(y)|^{2}$, for the two Zeeman-split modes in the $K$ valley at $E=0$: (a) spin up and (b) spin down. The sparse-stitching interface is well matched and so maximally transparent. The degree of $e$-$h$ hybridization is given in Table 2. The magenta line is the zero-Zeeman case for comparison. [Correspond to spectrum in Fig. 16(a).]
@@
Figure 18: For an imperfect interface, wavefunctions for the two $E\!=\!0$ modes in the $K$ valley as they are split by the Zeeman effect: (a) spin up and (b) spin down. Here, $\mu_{S}/t=4.2$. The magenta line is the zero-Zeeman case for comparison (not plotted in superconductor for clarity). [Correspond to spectrum in Fig. 16(b).]
@@ VII Conclusions
We have presented results for an infinite zigzag graphene/superconductor interface [77] using continuum and tight-binding models in the standard mean-field approach to spatially inhomogeneous superconductivity (BdG equation). Our main assumptions are: (i) disorder is absent and (ii) the fields $\Delta$, $B$, and $U$ are taken to be step functions at the interface. The continuum results are particularly helpful in exploring a wide range of parameters.
The primary conclusion is that strong Andreev reflection, and hence strong electron-hole hybridization, is much easier to achieve with graphene than with a typical semiconductor 2D electron gas. The reason is that the velocity at the Dirac points in graphene is large (of order the Fermi velocity of metals), even though the conduction electron density (and hence $k_{F}$) is low [71]. Furthermore, the Dirac point velocity is independent of the graphene chemical potential or density. Thus, current continuity across the interface can be satisfied without a great deal of electron reflection.
In order to actually realize strong Andreev reflection, the interface should be fairly transparent. In particular, there should be no obvious barrier, and the superconductor should be well-matched, having Fermi level states with wavevector larger than $K$ and velocity of order $v$ (the wave vector and velocity of the Dirac point). This does not appear to be a stringent restriction, since the Fermi surface of metals used for thin-film superconductors often have several sheets, allowing for several possible matches.
Previous work has emphasized that fine-tuning is necessary in order to achieve electron-hole hybridization at a clean QH/superconductor interface [45, 58, 59]; for instance, the chemical potential in the QH material should be tuned on an energy scale less than $E_{L}$. We emphasize that this is not the case for graphene. Once one obtains a fairly well-matched transparent interface, which may involve “coarse-tuning” the choice of superconductor on an energy scale of the bandwidth ($\gg\!E_{L}$), the $e$-$h$ hybridization in a graphene CAEM is ideal and independent of filling. Indeed, the dispersion of the CAEM self-aligns to be antisymmetric within each valley, independent of filling. Thus, the spectrum is valley degenerate, and the guiding center coordinate of the state at the chemical potential is at the graphene/superconductor interface.
@@
In the spatial profile of the wavefunction, the weight of the electron component is primarily on the sublattice corresponding to the bulk LLL state (sublattice-valley locking), and is surprisingly close to a truncated bulk wavefunction. This demonstrates that the terminated-lattice QH edge states are a poor starting point for understanding the CAEM of a transparent interface, an unsurprising result in hindsight in view of the very different boundary conditions. Rather, in the case of strong graphene-superconductor coupling ($t_{NS}\!\gg\!E_{L}\!\gg\!\Delta$), bulk-like electron and hole wavefunctions can be made at $E\!=0$ using $t_{NS}$, which are then coupled by superconductivity ($\Delta$).
In future work, the role of disorder, both in the superconductor and along the interface, should be further developed in the context of a transparent graphene/superconductor interface (for initial work see Ref. [54] and for the importance of disorder in the tunneling regime see Ref. [58]). We suspect that disorder will decrease the dependence on $\mu_{S}$ in our models and so increase the window of good electron-hole hybridization. A second clear direction is to include the penetration of the magnetic field into the superconductor in the form of vortices. It is likely that tunneling of electrons into states in the vortices is an important source of loss in the experiments [29, 30], and theoretical work assessing the importance of this effect has recently appeared [56, 61, 64]. Third, the role of spin-orbit coupling should be investigated. While such coupling is weak in graphene, in several of the superconductors typically used it is strong.
Finally, perhaps the most straight forward extension of our results is to the properties of nanostructures, which involve scattering between these CAEM at junctions, corners, and quantum point contacts. This would make contact with transport experiments, typically done in nanostructures [16, 17, 18, 29, 30]. A preliminary step toward the experiments can be made using the infinite-interface property $f_{e}$ calculated here [Eq. (1)]. Suppose the interface is connected to incoming and outgoing graphene terminated-lattice edge states. If one assumes that the electron and hole amplitudes on the outgoing edge directly reflect the hybridization of the interface states, then the probability that an electron exits in the outgoing edge state is simply $f_{e}$ and the downstream resistance, $R_{d}$, follows [29]:
$P_{e}=f_{e}\;,\quad R_{d}=R_{H}\frac{2f_{e}-1}{1-f_{e}}\;,$ (26)
where $R_{H}$ is the Hall resistance. However, the assumption is not usually valid: it is known that both valley momentum conservation [46] and scattering at corners [60] are important, even in the clean case. Having focused on properties of the infinite interface here, we leave consideration of these nanostructure effects to future work.
@@ Acknowledgements.
We thank Ethan Arnault, Anthony David, Gleb Finkelstein, William Klein, and Lingfei Zhao for helpful discussions. The work in the USA was supported in part by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Materials Sciences and Engineering Division, under Award No. DE-SC0005237. A.B. acknowledges support by the DoD National Defense Science and Engineering Graduate Fellowship Program. G.Z. acknowledges support from the National Natural Science Foundation of China (Grant No. 12374158) and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302400).
@@
Appendix A Specifics of the Models:
Tight-Binding and Continuum
@@ Appendix A Specifics of the Models:
$H_{T,\text{dense}}=$ $H_{T}^{(\textrm{A})}+H_{T}^{(\textrm{B})}$